OpenStudy (anonymous):

solve for x. -6x + 11 ≤ -9x + 2

OpenStudy (lgbasallote):

isolate x as if you're solving an equality.. \[-6x + 9x \le -11 + 2\] \[3x \le -9\] \[x \le -3\]